# TextCrafter: Accurately Rendering Multiple Texts in Complex Visual Scenes

### v3

@@ Abstract
Figure 1: TextCrafter enables precise multi-region visual text rendering, addressing the challenges of long, small-size, various numbers, symbols and styles in visual text generation. We illustrate the comparisons among TextCrafter with three state-of-the-art models, i.e., FLUX BlackForest [1], TextDiffuser-2 Chen et al. [5] and 3DIS Zhou et al. [45].
1 Introduction
@@
In recent years, diffusion models  [13, 27, 30, 28, 29, 40, 8, 6, 17] have emerged as the forefront technology in visual image generation. Recent studies, represented by general-purpose text-to-image models such as Flux [1] and SD 3.5 [9], have acquired the ability to render simple visual text through large-scale pretraining. However, when confronted with complex real-world visual text scenarios, these models often struggle with critical challenges such as text distortion, omission, and blurriness. These shortcomings significantly hinder their practical applicability in real-world use cases. Achieving precise, legible, and contextually accurate visual text rendering in complex scenes remains essential yet highly challenging.
@@
Recent studies in multi-instance control [10, 43, 7, 45, 18] primarily focus on generating multiple instances within a single image simultaneously, striving to achieve precise regulation over factors such as quantity, position and attributes. However, a crucial yet often overlooked category in the visual domain is visual text. Text is an indispensable component of the real world, distinguished by its unparalleled complexity and delicacy compared to conventional objects, flora, or fauna. Unlike these entities, text possesses an exceptionally intricate internal structure, where even minor alterations can cause significant changes in visual appearance, potentially leading to misrecognition or complete illegibility. Existing multi-instance approaches, such as MIGC [43] and 3DIS [45], primarily focus on sentence-level information and interaction, offering a limited understanding of text at a finer granularity. Moreover, RPG [35] and RAG-Diffusion [7] adjust instance structures during attention fusion but fall short of the precision and stability needed for visual text rendering.
@@
Recent studies on visual text rendering primarily focus on accurately generating a single text region. For instance, methods like AnyText [32], Glyph-byT5 [23], SceneTextGen [39], Diff-Text [38], and TextDiffuser [4] often employ specialized modules to encode visual text representations and condition the diffusion model accordingly. However, these methods face two major limitations: 1) They rely on predefined rules to synthesize visual text training datasets. Due to the need for highly accurate text annotations, manual verification is often required, making the creation of complex visual text datasets both labor-intensive and costly. 2) Most approaches depend on fine-tuned text encoders (e.g., Glyph-by-T5) or trained conditional control encoders (e.g., AnyText, Diff-Text, TextDiffuser, SceneTextGen) to facilitate text generation. In complex multi-text generation, however, these approaches often cause interference among the control features of different targets, making it difficult to balance fine-grained local control with global consistency across multiple texts.
@@
This paper explores the task of Complex Visual Text Generation (CVTG), which aims to generate intricate text content distributed across different regions. To tackle the challenges of CVTG, we propose TextCrafter, a novel training-free framework for rendering multiple texts in complex visual scenes. TextCrafter enhances multi-text control, ensuring accurate text presentation and logical layout. Our framework comprises three key stages: 1) Instance Fusion: This stage strengthens the connection between visual text and its corresponding carrier, ensuring strong consistency between the text content and its surrounding environment. It effectively prevents text from appearing in incorrect positions. 2) Region Insulation: Leveraging the positional priors of the pre-trained DiT model [26] (e.g., FLUX), this step initializes the layout information for each text instance while separating and denoising text prompts across different regions. This process prevents early interference between text areas, reducing confusion and the risk of content omission in multi-text scenarios. 3) Text Focus: An attention control mechanism is introduced to enhance the attention maps of visual text, refining the fidelity of text rendering. This stage ensures textual accuracy and addresses issues of blurriness, particularly in smaller text. We present some examples in Figure 1. Our contributions are summarized as follows:
@@
  • We propose TextCrafter, a novel training-free framework for rendering multiple texts in complex visual scenes, consisting of three key stages: instance fusion, region insulation, and token focus.
@@
  • We conduct extensive quantitative and qualitative experiments on the CVTG-2K benchmark. The results demonstrate that TextCrafter delivers exceptional performance, proving its superior effectiveness and robustness in tackling the CVTG task.
@@
2 Related Works
Multi-instance Compositional Generation.
Attend-and-Excite [3] and BOX-Diffusion [34] guide diffusion models to align with prompts. GLIGEN [21], MIGC [43], and CreativeLayout [37] incorporate bounding boxes into trainable layers. RPG [35], RAG-Diffusion [7] and DreamRenderer [44] decompose generation into regional tasks. However, these methods overlook visual text, a critical element for image realism.
@@ 2 Related Works
DiffSTE [19], Glyph-ByT5 [23], and UDiffText [41] use character-level encoders for word appearance. TextDiffuser [4] and GlyphControl [36] generate text layout masks. Diff-Text [38] restricts attention maps. GlyphDraw [24] utilizes dual models for positioning and rendering. TextDiffuser-2 [5] leverages LLMs for layout planning. These approaches optimize for single region or individual object, limiting generalization in complex scenes.
Benchmark for Visual Text.
GlyphControl introduced SimpleBench and CreativeBench [36], but these contain only single-word scenes. MARIO-Eval [4] has low-quality prompts. DrawTextExt Creative [24] lacks sufficient coverage. AnyText-benchmark [32] treats words separately, limiting distribution. Existing benchmarks focus on a single object or location, failing to capture real-world complexity. Thus, we propose CVTG-2K to evaluate models with diverse positions, quantities, lengths, and attributes in complex visual text scenarios.
Figure 3: Illustration of tokenizing the prompt “A sidewalk poster with ‘Register Now for IJCAI 2025’.” along with the attention map corresponding to each token. The use of preceding quotation marks can reinforce the relationship between text tokens and carrier tokens.
Figure 4: Motivation for using a layout optimizer. LLM/MLLM-based layout designers frequently generate overlapping bounding boxes, which do not align with the layout preferences of Diffusion models. Our proposed layout optimizer addresses this limitation without requiring additional fine-tuning.
3 Methodology
3.1 Overview
@@
In CVTG, users provide a global prompt $P$ containing multiple visual text descriptions $D=\{d_{1},d_{2},\dots,d_{n}\}$, where each description includes the text’s main body and descriptors (position, attributes). The main text bodies are defined as $VT=\{vt_{1},vt_{2},\dots,vt_{n}\}$, with each $vt_{i}$ corresponding to descriptors $d_{i}$. The model must generate an image from prompt $P$ where each $vt_{i}$ appears with its corresponding description $d_{i}$.
@@
Current text-to-image models excel with individual visual texts. By decomposing CVTG into simpler sub-tasks, overall complexity decreases. Solving these sub-tasks separately and combining their solutions accomplishes the CVTG task. This approach shifts focus progressively from the entire task to sub-tasks then to specific visual texts, enabling precise rendering. Based on this coarse-to-fine strategy, we present TextCrafter (Figure 2), detailed in subsequent sections.
3.2 Instance Fusion
It is crucial to prevent visual texts from floating in the image and ensure proper positioning. We propose an instance fusion strategy that integrates the visual text’s content with its spatial carrier. Rather than using the text’s own embedding, we leverage embedding of the preceding quotation mark to establish spatial relationships. As illustrated in Figure 3, the preceding quotation mark in the attention map corresponds to the visual text it governs, indicating that it encapsulates complete information about the text. According to token additivity [14], we incorporate the quotation mark’s embedding into the carrier’s embedding via weighted fusion to ensure precise alignment, where $\lambda$ controls the proportion of the addition.
@@ 3.2 Instance Fusion
Table 1: Comparison of CVTG-2K with existing public visual text Benchmarks in terms of Number of Samples (Number), Average Word Count (Word), Average Character Count (Char), Attribute, and number of Regions (Region).
Benchmark Number Word Char Attribute Region
@@
AnyText-benchmark 1000 4.18 21.84 $\times$ $\times$
@@
size/color/font
@@
1:  Input: The encoded embedding of a global prompt $P$ is denoted as $c$. The encoded embedding of multiple descriptions $D$ is denoted as $\{\hat{c}^{1},\hat{c}^{2},\dots,\hat{c}^{n}\}$. 2:  Parameter: the proportion of the addition $\lambda$, the $r$ steps of Region Insulation, the pre-generation steps $\tau$. 3:  Output: A generated latent $x_{0}$ 4:  {Step 1: Instance Fusion} 5:  $c=\text{InstanceFusion}(c,\lambda)$ 6:  {Step 2: Region Insulation} 7:  $z_{t}\sim\mathcal{N}(0,1)$ 8:  for $t=\tau,\tau-1,\dots,1$ do 9:     $z_{t-1},A_{t}\leftarrow\text{DiT}(z_{t},c,t)$ 10:  end for 11:  for $i=1,2,\dots,n$ do 12:     $p_{i}\leftarrow(0,0)$ 13:     $p_{\text{max},i}\leftarrow\arg\max_{p_{i}}A_{t}(p_{i},i)$ 14:  end for 15:  $bbx_{1},\dots,bbx_{n}\leftarrow\text{MILP}(p_{\text{max},1},\dots,p_{\text{max},n})$ 16:  $x_{T}\sim\mathcal{N}(0,1)$ 17:  for $i=1,2,\dots,n$ do 18:     $\hat{x}_{t}^{i}\leftarrow\text{Initialize}(\mathcal{N}(0,1),bbx_{i})$ 19:  end for 20:  for $t=T,T-1,\dots,T-r+1$ do 21:     $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$ 22:     for $i=1,2,\dots,n$ do 23:        $\hat{x}_{t-1}^{i}\leftarrow\hat{x}_{t}^{i}-\epsilon_{\theta}(\hat{x}_{t}^{i},\hat{c}^{i})$ 24:        $x_{t-1}\leftarrow\text{Reinsert}(x_{t-1},\hat{x}_{t-1}^{i},bbx_{i})$ 25:     end for 26:  end for 27:  {Step 3: Text Focus} 28:  for $t=T-r,T-r-1,\dots,1$ do 29:     for $\text{AttentionLayer}_{i}\in\epsilon_{\theta}$ do 30:        $M_{t}=Q_{t}^{i}\times K_{t}^{i}$ 31:        $M_{t}\leftarrow\text{Focus}(M_{t})$ 32:        $\text{Attention}_{t}^{i}\leftarrow\text{softmax}(M_{t}/\sqrt{d})\times V_{i}$ 33:     end for 34:     $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$ 35:  end for 36:  return  $x_{0}$
Table 2: Quantitative results on the proposed CVTG-2K. The results highlight TextCrafter’s superiority across five metrics.
@@
FLUX.1 dev (2024) 0.4965 0.6879 0.7401 0.8886 5.91
SD3.5 Large (ICML 2024) 0.6548 0.8470 0.7797 0.9297 5.56
AnyText (ICLR 2024) 0.1804 0.4675 0.7432 0.6935 4.53
TextDiffuser-2 (ECCV 2024) 0.2326 0.4353 0.6765 0.5627 4.51
3DIS (ICLR 2025) 0.3813 0.6505 0.7767 0.8684 4.86
@@
3.3 Region Insulation
@@
In CVTG, multiple texts often influence each other’s glyphs and layout [42]. To simplify, we decouple them into independent instances. Specifically, we assign a rectangular bounding box to each visual text to represent its layout, initialized based on the positional preferences of the pre-trained DiT model. This allows us to effectively isolate the texts, preventing interference during generation. Instead of manually specifying layouts or relying on LLMs, we leverage the diffusion model’s learned positional knowledge [31, 16]. Figure 4 demonstrates a notable limitation of LLMs/MLLMs: despite dedicated fine-tuning, the layouts they provide often exhibit overlapping bounding boxes and do not conform to the layout preferences inherent in Diffusion models. As shown in Figure 5, during the early denoising steps, we capture attention maps and identify the points of maximum attention for each visual text. This pre-generation process guides layout initialization. Assume that the attention map obtained during this process is $A_{t}$, where $t$ represents the current step, and $p$ stands for pixel:
@@
Next, the rectangular bounding boxes for each visual text are optimized using Mixed-Integer Linear Programming (MILP) to minimize the Manhattan distance between the bounding box center and the attention point. Let $p_{\text{max},i}$ be the maximum attention point of $vt_{i}$ and $c_{i}$ the rectangular bounding box center. The optimization objective is:
$\min\sum_{i}D_{\text{Manhattan}}(p_{\text{max},i},c_{i}).$ (2)
Simultaneously incorporate a set of constraints to ensure that the layout of each visual text is reasonable and non-overlapping. The solution to this optimization problem can be represented as assigning a bounding box $bbx_{i}$ $(m_{\text{offset}},n_{\text{offset}},m_{\text{scale}},n_{\text{scale}})$ to each description. Then, we encode the prompt $P$ and the corresponding descriptions of visual text $D=\{d_{1},d_{2},\dots,d_{n}\}$, obtaining values for $c$ and $\{\hat{c}^{1},\hat{c}^{2},\dots,\hat{c}^{n}\}$. $c$ needs to be processed through instance fusion introduced in Section 3.2. The latent $\hat{x}_{i}$ is conditioned on $\hat{c}_{i}$, and the initial latent $x$ is conditioned on the complete $c$. When $t>T-r$ (i.e., during the first $r$ steps of the denoising process), denoising follows the formula:
@@
3.4 Text Focus
Region insulation creates distinct visual boundaries between visual texts. To enhance harmony, regions are merged into a single latent variable for denoising, but this may blur smaller texts. To preserve structure, we boost attention scores [11, 15] for texts and preceding quotation marks, capping enhancement at 2$\times$ to prevent overamplification. To ensure control, we employ the $\tanh$ function. The $k$ token sequence is denoted as $F=\{f_{1},\dots,f_{k}\}$. The enhancement ratio is:
@@ 3.4 Text Focus
We use a smaller ratio for single-word cases. Unlike UNet’s cross-attention, MM-DiT employs a full attention matrix with four regions: image-to-image, prompt-to-prompt, prompt-to-image, and image-to-prompt [2] . We re-weight the image-to-text matrix, which controls image modality and enhances focus on visual text. At step $t$, the operation for each element in $M_{t}$ is expressed as:
@@
The overall algorithmic process is illustrated in Algorithm 1.
Table 3: Ablation studies on CVTG-2K dataset, including Instance Fusion (Step 1), Region Insulation (Step 2), Text Focus (Step 3), and Layout methods.
Steps Layout Method
Metrics S1 S2 S3 S1+S2 S2+S3 All Random Fixed Optimizer
Word Accuracy↑ 0.4422 0.6116 0.6351 0.6121 0.7328 0.7370 0.6900 0.6994 0.7370
NED↑ 0.6495 0.7750 0.8055 0.7762 0.8651 0.8679 0.8374 0.8475 0.8679
@@
4 Experiments
4.1 Benchmark
Currently, there is a lack of publicly available benchmark datasets specifically designed for complex visual text generation tasks. Therefore, we propose CVTG-2K, a challenging public benchmark dataset for complex visual text. All prompts are generated through the OpenAI’s O1-mini API [25]. These prompts encompass a variety of scenes containing complex visual texts, including but not limited to street views, book covers, advertisements, posters, notes, memes, logos, and movie frames. Unlike previous standard datasets synthesized using fixed rules, CVTG-2K ensures the diversity and rationality of the data distribution. We first prompt the O1 model to conceive a scene and then imagine the spossible visual texts that may appear in that scene, utilizing Chain-of-Thought (CoT) [33] techniques to ensure the quality of the generated prompts. For further details on the dialogue template with O1, refer to the supplementary material. Through rigorous filtering rules and meticulous post-processing, we have generated 2,000 prompts containing complex visual texts. On average, the visual texts in CVTG-2K contain 8.10 words and 39.47 characters, surpassing all previously published visual text benchmark datasets in terms of visual text length. Table 1 presents a comparison of the data in CVTG-2K with currently available benchmark datasets. Furthermore, CVTG-2K is the first benchmark dataset to include the number of multiple visual text regions, with the number of regions ranging from 2 to 5. We have assigned different proportions to these region numbers, approximately 20%, 30%, 30%, and 20%. With its diverse region numbers and visual text lengths, CVTG-2K provides a comprehensive evaluation of model performance in complex visual text generation tasks. Through careful human review and filtering, CVTG-2K ensures that it does not contain any discriminatory or inflammatory content.

### v6

@@ Abstract
Figure 1: TextCrafter enables precise multi-region visual text rendering, addressing the challenges of long, small-size, various numbers, symbols and styles in visual text generation. We illustrate the comparisons among TextCrafter with three state-of-the-art models, i.e., FLUX (BlackForest 2024), TextDiffuser-2 (Chen et al. 2025) and 3DIS (Zhou et al. 2025b).
Introduction
@@
In recent years, diffusion models  (Ho, Jain, and Abbeel 2020; Podell et al. 2023; Rombach et al. 2022; Ramesh et al. 2021, 2022; Zhao et al. 2024; Dhariwal and Nichol 2021; Chen et al. 2023) have emerged as the forefront technology in image generation. Recent studies, represented by general-purpose text-to-image models such as FLUX (BlackForest 2024) and SD 3 (Esser et al. 2024), have acquired the ability to render simple visual text through large-scale pretraining. However, when confronted with complex real-world visual text scenarios, these models often struggle with critical challenges such as text distortion, omission, and blurriness. These shortcomings significantly hinder their practical applicability in real-world use cases. Achieving precise, legible, and contextually accurate visual text rendering in complex scenes remains essential yet highly challenging.
@@
Recent studies in multi-instance generaion (Feng et al. 2022; Zhou et al. 2024; Chen et al. 2024b; Zhou et al. 2025b) primarily focus on generating multiple instances within a single image simultaneously, striving to achieve precise regulation over factors such as quantity, position and attributes. However, a crucial yet often overlooked category in the visual domain is visual text. Text is an indispensable component of the real world, distinguished by its unparalleled complexity and delicacy compared to conventional objects, flora, or fauna. Unlike these entities, text possesses an exceptionally intricate internal structure, where even minor alterations can cause significant changes in visual appearance, potentially leading to misrecognition or complete illegibility. Existing multi-instance generaion approaches, such as MIGC (Zhou et al. 2024) and 3DIS (Zhou et al. 2025b), primarily focus on sentence-level information and interaction, offering a limited understanding of text at a finer granularity. Moreover, RPG (Yang et al. 2024a) and RAG-Diffusion (Chen et al. 2024b) adjust instance structures during attention fusion but fall short of the precision and stability needed for visual text rendering.
@@
Recent studies on visual text rendering primarily focus on accurately generating a single text region. For instance, methods like AnyText (Tuo et al. 2024), Glyph-byT5 (Liu et al. 2025), SceneTextGen (Zhangli et al. 2024), Diff-Text (Zhang et al. 2024b), and TextDiffuser (Chen et al. 2024a) often employ specialized modules to encode visual text representations and condition the diffusion model accordingly. However, these methods face two major limitations: 1) They rely on predefined rules to synthesize visual text training datasets. Due to the need for highly accurate text annotations, manual verification is often required, making the creation of complex visual text datasets both labor-intensive and costly. 2) Most approaches depend on fine-tuned text encoders (e.g., Glyph-by-T5) or trained conditional control encoders (e.g., AnyText, TextDiffuser, SceneTextGen) to facilitate text generation. In complex multi-text generation, however, these approaches often cause interference among the control features of different targets, making it difficult to balance fine-grained local control with global consistency across multiple texts.
@@
This paper explores the task of Complex Visual Text Generation (CVTG), which aims to generate intricate text content distributed across different regions. To tackle the challenges of CVTG, we propose TextCrafter, a novel training-free framework for rendering multiple texts in complex visual scenes. TextCrafter enhances multi-text control, ensuring accurate text presentation and logical layout. Our framework comprises three key stages: 1) Instance Fusion: This stage strengthens the connection between visual text and its corresponding carrier, ensuring strong consistency between the text content and its surrounding environment. It effectively prevents text from appearing in incorrect positions. 2) Region Insulation: Leveraging the positional priors of the pre-trained DiT model (Peebles and Xie 2023), this step initializes the layout information for each text instance while separating and denoising text prompts across different regions. This process prevents early interference between text areas, reducing confusion and the risk of content omission in multi-text scenarios. 3) Text Focus: An attention control mechanism is introduced to enhance the attention maps of visual text, refining the fidelity of text rendering. This stage ensures textual accuracy and addresses issues of blurriness, particularly in smaller text. We present some examples in Figure 1. Our contributions are summarized as follows:
@@
  • We propose TextCrafter, a novel training-free framework for rendering multiple texts in complex visual scenes, consisting of three key stages: Instance Fusion, Region Insulation, and Text Focus.
@@
  • We conduct extensive quantitative and qualitative experiments on the CVTG-2K and a established benchmark. The results demonstrate that TextCrafter exhibits superior effectiveness and robustness in the CVTG task.
@@
Related Works
Compositional Generation and Multi-instance Generation.
Attend-and-Excite (Chefer et al. 2023) and BOX-Diffusion (Xie et al. 2023) guide pre-trained diffusion models to align generated instances with prompts. GLIGEN (Li et al. 2023), MIGC (Zhou et al. 2024), and CreativeLayout (Zhang et al. 2024a) enhance control by incorporating bounding box conditions into trainable layers. RPG (Yang et al. 2024a), RAG-Diffusion (Chen et al. 2024b) and DreamRenderer (Zhou et al. 2025a) break the generation process into regional tasks for compositional generation. However, these methods often overlook a critical kind of instance: visual text, as a key component of visual scenes.
@@
DiffSTE (Ji et al. 2023), Glyph-ByT5 (Liu et al. 2025), and UDiffText (Zhao and Lian 2025) use character-level encoders to incorporate word appearance into embeddings. TextDiffuser (Chen et al. 2024a) creates text layout masks for injection into the latent space. Approaches like Diff-Text (Zhang et al. 2024b) improve text accuracy with attention map restrictions. GlyphDraw (Ma et al. 2023) uses two models for text position prediction and rendering. TextDiffuser-2 (Chen et al. 2025) leverages large language models for layout planning. However, these approaches often optimize text rendering for single region, leading to poor generalization in complex visual text scenes.
Benchmark for Visual Text Generation.
Earlier works like GlyphControl introduced SimpleBench and CreativeBench (Yang et al. 2024b), but these fixed-template datasets offer limited diversity and contain only single-word scenes. MARIO-Eval (Chen et al. 2024a) suffers from low-quality prompts and unclear semantics. Although DrawTextExt (Ma et al. 2023) collects prompts for generating natural scenes, the comprehensiveness and scale of the dataset are insufficient. AnyText-benchmark (Tuo et al. 2024) allows multi-word prompts but treats each word separately and appends them directly to the caption, limiting data distribution. Existing visual text benchmarks focus on text generation within a single object or location, falling short of capturing the complexity of real-world scenes. Thus, we propose a benchmark CVTG-2K to incorporate diverse positions, quantities, lengths, and attributes for comprehensively evaluate models in complex visual text scenarios.
Figure 3: Illustration of tokenizing the prompt “A sidewalk poster with ‘Register Now for arXiv’.” along with the attention map corresponding to each token. The use of preceding quotation marks can reinforce the relationship between text tokens and carrier tokens.
Figure 4: Motivation for layout optimizer. LLM/MLLM-based layout designers frequently generate overlapping bounding boxes, which do not align with the layout preferences of Diffusion models. Our proposed layout optimizer addresses this limitation without additional fine-tuning.
Methodology
Overview
@@
In CVTG, users provide a global prompt $P$ containing multiple visual text descriptions $D=\{d_{1},d_{2},\dots,d_{n}\}$, where each description includes the visual text’s content and descriptors (position, attributes). The visual texts’ content are defined as $VT=\{vt_{1},vt_{2},\dots,vt_{n}\}$, with each $vt_{i}$ corresponding to description $d_{i}$. The model must generate an image from prompt $P$ where each $vt_{i}$ appears with its corresponding description $d_{i}$.
@@
Current text-to-image models excel with individual visual texts. By decomposing CVTG into simpler sub-tasks, overall complexity decreases. Solving these sub-tasks separately and combining their solutions accomplishes the CVTG task. This approach shifts focus progressively from the entire task to sub-tasks then to specific visual texts, enabling precise rendering. Based on this coarse-to-fine strategy, we present TextCrafter (Figure 2).
Novelty.
TextCrafter differs significantly from existing compositional generation and multi-instance generation (collectively referred to as MIC) methods. MIC methods excel at generating multiple instances of different categories, but often suffer from feature interference when handling highly similar instances, leading to corruption of internal instance structure. TextCrafter addresses this issue through a dual mechanism: while MIC methods treat each instance as an independent entity during generation, TextCrafter employs Instance Fusion to bind text instances with their visual carriers, subsequently treating them as unified wholes during generation, thereby enhancing distinctions between different text instances. Additionally, the MILP layout optimizer generates non-overlapping layouts through hard constraints, eliminating overlapping regions. Overlapping or mutually contained regions are extremely common in MIC tasks, highlighting an inherent difference between MIC and CVTG tasks. By separating the latent spaces of text instances from the initial generation stage, Region Insulation fundamentally prevents feature leakage caused by content similarity.
Instance Fusion
It is crucial to prevent visual texts from floating on the surface and ensure proper positioning. We propose an Instance Fusion strategy that integrates the visual text’s content with its spatial carrier. Rather than using the text’s own embedding, we leverage embedding of the preceding quotation mark to establish spatial relationships. As illustrated in Figure 3, the preceding quotation mark in the attention map corresponds to the visual text it governs, indicating that it encapsulates complete information about the text. According to token additivity (Hu et al. 2024), we incorporate the preceding quotation mark’s embedding into the carrier’s embedding via weighted fusion to ensure precise alignment, where $\lambda$ controls the proportion of the addition.
@@
Benchmark Num Word Char Attribute Region
@@
AnyText-
benchmark 1000 4.18 21.84 $\times$ $\times$
@@
size/color/
font
@@
Table 1: Comparison of CVTG-2K with existing public visual text Benchmarks in terms of Number of Samples (Num), Average Word Count (Word), Average Character Count (Char), Attribute, and number of Regions (Region).
@@
1:Input: The encoded embedding of a global prompt $P$ is denoted as $c$. The encoded embedding of multiple descriptions $D$ is denoted as $\{\hat{c}^{1},\hat{c}^{2},\dots,\hat{c}^{n}\}$.
2:Parameter: the proportion of the addition $\lambda$, the $r$ steps of Region Insulation, the pre-generation steps $\tau$.
3:Output: A generated latent $x_{0}$
4: {Step 1: Instance Fusion}
[…]
6: {Step 2: Region Insulation}
[…]
8:for $t=\tau,\tau-1,\dots,1$ do
9:  $z_{t-1},A_{t}\leftarrow\text{DiT}(z_{t},c,t)$
10:end for
11:for $i=1,2,\dots,n$ do
12:  $p_{i}\leftarrow(0,0)$
13:  $p_{\text{max},i}\leftarrow\arg\max_{p_{i}}A_{t}(p_{i},i)$
14:end for
[…]
17:for $i=1,2,\dots,n$ do
18:  $\hat{x}_{t}^{i}\leftarrow\text{Initialize}(\mathcal{N}(0,1),bbx_{i})$
19:end for
20:for $t=T,T-1,\dots,T-r+1$ do
21:  $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$
22:  for $i=1,2,\dots,n$ do
23:   $\hat{x}_{t-1}^{i}\leftarrow\hat{x}_{t}^{i}-\epsilon_{\theta}(\hat{x}_{t}^{i},\hat{c}^{i})$
24:   $x_{t-1}\leftarrow\text{Reinsert}(x_{t-1},\hat{x}_{t-1}^{i},bbx_{i})$
25:  end for
26:end for
27: {Step 3: Text Focus}
28:for $t=T-r,T-r-1,\dots,1$ do
29:  for $\text{AttentionLayer}_{i}\in\epsilon_{\theta}$ do
30:   $M_{t}=Q_{t}^{i}\times K_{t}^{i}$
31:   $M_{t}\leftarrow\text{Focus}(M_{t})$
32:   $\text{Attention}_{t}^{i}\leftarrow\text{softmax}(M_{t}/\sqrt{d})\times V_{i}$
33:  end for
34:  $x_{t-1}\leftarrow x_{t}-\epsilon_{\theta}(x_{t},c)$
35:end for
36:return  $x_{0}$
@@
FLUX.1 dev [2024] 0.4965 0.6879 0.7401 0.8886 5.91
SD3.5 Large [ICML 2024] 0.6548 0.8470 0.7797 0.9297 5.56
AnyText [ICLR 2024] 0.1804 0.4675 0.7432 0.6935 4.53
TextDiffuser-2 [ECCV 2024] 0.2326 0.4353 0.6765 0.5627 4.51
RAG-Diffusion [ICCV 2025] 0.2648 0.4498 0.6688 0.6397 5.58
3DIS [ICLR 2025] 0.3813 0.6505 0.7767 0.8684 4.86
@@
Table 2: Quantitative results on CVTG-2K dataset. Bold values denote the best performance, while underlined values indicate the second-best performance for each metric.
Region Insulation
@@
In CVTG, multiple texts often influence each other’s glyphs and layout (Zheng et al. 2023). To simplify, we decouple them into independent instances. Specifically, we assign a rectangular bounding box to each visual text to represent its layout. This allows us to effectively isolate the texts, preventing interference during generation. Instead of manually specifying layouts or relying on MLLMs, we leverage the diffusion model’s learned positional knowledge (Tumanyan et al. 2023). Figure 4 demonstrates a notable limitation of MLLMs: despite dedicated fine-tuning, the layouts they provide often exhibit overlapping bounding boxes and do not conform to the layout preferences inherent in Diffusion models. As shown in Figure 5, during the early denoising steps, we capture attention maps and identify the points of maximum attention for each visual text. This pre-generation process guides layout initialization. Assume that the attention map obtained during this process is $A_{t}$, where $t$ represents the current step, and $p$ stands for pixel:
@@
Next, the rectangular bounding boxes for each visual text are optimized using Mixed-Integer Linear Programming (MILP) to minimize the Manhattan distance (L1 norm) between the bounding box center and the maximum attention point. Let $p_{\text{max},i}$ be the maximum attention point of $vt_{i}$ and $p_{\text{center},i}$ the rectangular bounding box center. Thus, the optimization objective is:
$\min\sum_{i=1}^{n}\left|p_{\text{center},i}-p_{\text{max},i}\right|_{1}$ (2)
Additionally, we enforce constraints on the minimum area, aspect ratio, and non-overlapping properties of the bounding boxes. Further details of the MILP optimizer can be found in the Supplementary Material. The solution to this optimization problem can be represented as assigning a bounding box $bbx_{i}$ $(m_{\text{offset}},n_{\text{offset}},m_{\text{scale}},n_{\text{scale}})$ to each description $d_{i}$. Then, we encode the prompt $P$ and the corresponding descriptions of visual text $D=\{d_{1},d_{2},\dots,d_{n}\}$, obtaining values for $c$ and $\{\hat{c}^{1},\hat{c}^{2},\dots,\hat{c}^{n}\}$. $c$ needs to be processed through Instance Fusion. The latent $\hat{x}_{i}$ is conditioned on $\hat{c}_{i}$, and the initial latent $x$ is conditioned on the complete $c$. When t>Trt>T-r (i.e., during the first $r$ steps of the denoising process), denoising follows the formula:
@@
Text Focus
[…]
@@
We use a smaller ratio for single-word cases. Unlike UNet’s cross-attention, MM-DiT employs a full attention matrix with four regions: image-to-image, text-to-text, text-to-image, and image-to-text (Cai et al. 2024) . We re-weight the image-to-text matrix, which controls image modality and enhances focus on visual text. At step $t$, the operation for each element in $M_{t}$ is expressed as:
@@
The overall algorithm is illustrated in Algorithm 1.
@@
Experiments
Benchmark
CVTG-2K.
Currently, there is a lack of publicly available benchmark datasets specifically designed for complex visual text generation tasks. Therefore, we propose CVTG-2K, a challenging public benchmark dataset for complex visual text. All prompts are generated through the OpenAI’s O1-mini API (OpenAI 2024). These prompts encompass a variety of scenes containing complex visual texts, including but not limited to street views, book covers, advertisements, posters, notes, memes, logos, and movie frames. Unlike previous standard datasets synthesized using fixed rules, CVTG-2K ensures the diversity and rationality of the data distribution. We first prompt the O1 model to conceive a scene and then imagine the possible visual texts that may appear in that scene, utilizing Chain-of-Thought (CoT) (Wei et al. 2022) techniques to ensure the quality of the generated prompts. For further details on the dialogue template with O1, refer to the supplementary material. Through rigorous filtering rules and meticulous post-processing, we have generated 2,000 prompts containing complex visual texts. On average, the visual texts in CVTG-2K contain 8.10 words and 39.47 characters, surpassing all previously published visual text benchmark datasets in terms of visual text length. Table 1 presents a comparison of the data in CVTG-2K with currently available benchmark datasets. Furthermore, CVTG-2K is the first benchmark dataset to include the number of multiple visual text regions, with the number of regions ranging from 2 to 5. We have assigned different proportions to these region numbers, approximately 20%, 30%, 30%, and 20%. With its diverse region numbers and visual text lengths, CVTG-2K provides a comprehensive evaluation of model performance in complex visual text generation. Through careful review, CVTG-2K ensures that it does not contain any discriminatory or inflammatory content.